----------------------------------------------------------------------------------------------- 8:00 keep taking hydrogen elastic data at 1 uA. 8:47 restart coda because end of run information was lost for run 3367. 9:40 move target to 3He. raster size is 2X2 10:00 We change the half wave plate to out and start to take 3He production data. 10:11 We make the coda download again, because Bob changed the trigger supervisor code. 10:18 we change the prescale since we change the setting to 2uA new prescale are 20000:100:1 10:35 beam is gone because MCC is going to check the laser. 11:16 changed the BB shower threshold to -145 mV 12:15 MCC fixed the laser problem, beam is back 12:16 check the raster size. 12:50 the BB HV tripped when we try to go 5uA from 2uA 12:50 the beam current goes to 5uA. 13:30 MCC test plan NMR measurement and target goes to empty. 14:30 MCC test plan break, take production data.
------------------------------------------------------------------------------------------------
Run 3367 7:40am-8:14 H2 target 1uA events 3M
Run 3368 8:14-8:47 H2 target 1uA events 3M
Run 3369 8:50-9:25 H2 target 1uA events 3M
Run 3370 9:25-9:40 H2 target 1uA events 1.1M
Run 3371 9:58-9:59 3He target 2uA to check raster size
Run 3372 9:59-10:10 3He target 2uA production events events 2.1M
Run 3373 10:11-10:17 3He target 2uA events 1.4M prescale wrong
Run 3374 10:18-10:35 3He target 2uA events 0.67M right prescale
Run 3375 12:12-12:46 3He target 2uA events 0.28M right prescale
Run 3376 12:50-13:30 3He target 5uA events 0.97M right prescale
(20000,100,1)
Run 3377 14:30- 3He target 5uA events right prescale